OpenStudy (anonymous):

For the equation given below, evaluate dy/dx at the point (-1,2). (3x^2)-(3xy)+(y^3)=17

OpenStudy (anonymous):

i just did this problem in my math homework haha

OpenStudy (anonymous):

hahaha o gosh webwork?!

OpenStudy (anonymous):

so i just realized i did it wrong in my stuff haha. im gonna try and redo it now lol

OpenStudy (baldymcgee6):

Did you get it or do you still need help?

OpenStudy (anonymous):

o i still need help! thank you!

OpenStudy (baldymcgee6):

Okay, so we have to differentiate this implicitly, do you know what that is?

OpenStudy (anonymous):

yes! i get -4x+3y/-3x+3y^2

OpenStudy (anonymous):

but i don't get it from there....!

OpenStudy (baldymcgee6):

I think you're derivative is wrong...

OpenStudy (baldymcgee6):

your*

OpenStudy (anonymous):

i dont get it! please help

OpenStudy (baldymcgee6):

okay... here we go

OpenStudy (anonymous):

thank you!

OpenStudy (baldymcgee6):

http://screencast.com/t/x5qFeqZYY

OpenStudy (anonymous):

does that mean i can just plug in (-1,2) for x and y? because i get 4/3 and it says thats not correct!

OpenStudy (baldymcgee6):

hmm, you should just be able to plug those in...

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{2 x-y}{x-y^2} \]\[\frac{4}{5} \text{at } (-1,2) \]

OpenStudy (anonymous):

o thank you @robtobey! do u mind writing out the steps, im having trouble as to how u arrived to that. thanks so much!

OpenStudy (baldymcgee6):

@tenistaego my original solution just missed a y^2 in the denominator, this should be correct and equivalent to robtoobey's

OpenStudy (anonymous):

thx@baldymcgee6!

OpenStudy (baldymcgee6):

No problem, please choose the appropriate best answer :)

OpenStudy (anonymous):

A solution using Mathematica for the calculations is attached.

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

Your are welcome.